when one-third of a number is added to one-fourth of the number, the sum is 21. what is the number?
Answer by stanbon(75887)

when one-third of a number is added to one-fourth of the number, the sum is 21. what is the number?
------
Let the number be "x":
(1/3 + 1/4)x = 21
[(4+3)/12]x = 21
------
(7/12)x = 21
x = 12*21/7
x = 36
============
Cheers,
Stan H.